FORECAST ACTUAL
1,470 1,520
1,370 1,470
1,670 1,570
1,720 1,620
1,770 1,670

a. Compute the tracking signal using the mean absolute deviation and running sum of forecast errors. (Negative values should be indicated by a minus sign. Round your "Mean Absolute Deviation", "Tracking Signal" to 2 decimal places and all other answers to the nearest whole number.)


Period Forecast Actual Deviation RSFE Absolute
Deviation
Sum of
Absolute
Deviation
MAD TS
1 1,470 1,520
2 1,370 1,470
3 1,670 1,570
4 1,720 1,620
5 1,770 1,670
Answer #1

a.We have,

Deviation=Actual -Forecast

Deviation in period 1=B3-C3=50

(Drag this cell formula for rest of the periods)

Running sum of forecast errors=RSFE of period 1=Deviation of period 1 =50

RSFE of period 2 =E3+D4=150

(Drag this cell formula for rest of the periods)

Absolute deviation =Absolute value of (deviation)

Sum of absolute deviation of period 1=absolute deviation of period 1=50

Sum of absolute deviation of period 2=G3+F4=150

(Drag this cell formula for rest of the periods)

MAD of period 1=Absolute deviation of period 1=50

MAD of period 2=SUM($F$3:F4)/A4=75

(Drag this cell formula for rest of the periods)

Tracking signal =RSFE/MAD

Tracking signal of period 1=E3/H3=1

(Drag this cell formula for rest of the periods)